OpenStudy (anonymous):

solve for x (x-2)/(x+5) = (x+2)/(x+3)

OpenStudy (anonymous):

cross multiply and post what you get

OpenStudy (anonymous):

if I cross multiply then I get 6/10 or 2/5

OpenStudy (anonymous):

but I am pretty sure you don't just cross multiply

OpenStudy (anonymous):

not quite

OpenStudy (anonymous):

actually you do cross multiply

OpenStudy (anonymous):

exactly my point. I don't understand so I need help. I am sure you do, but there has to be more too it.

OpenStudy (saifoo.khan):

(x-2) (x+2) ----- = --------- (x+5) (x+3)

OpenStudy (saifoo.khan):

Cross multiply.

OpenStudy (anonymous):

Cross multiply is voodoo, but I suppose it helps some people remember

OpenStudy (anonymous):

okay, well this is what we will do: (x-2)(x+3)=x^2+x-6 (x+2)(x+5)=x^2+7x+10

OpenStudy (anonymous):

i did cross multiply and got the wrong answer

OpenStudy (anonymous):

vjay, stop that.

OpenStudy (anonymous):

now: x^2+x-6=x^2+7x+10

OpenStudy (anonymous):

now subtract x^2 from both sides

OpenStudy (saifoo.khan):

(x-2) (x+3) = (x+2)(x+5) x^2 +3x - 2x - 6 = x^2 + 5x + 2x + 10

OpenStudy (anonymous):

subtract 10 from both sides

OpenStudy (anonymous):

then subtract x from both sides

OpenStudy (anonymous):

you should be left with : -16=6x

OpenStudy (anonymous):

divide both sides by 6 and get: -16/6=x, which simplifies to -8/3 =x

OpenStudy (anonymous):

satified?

OpenStudy (saifoo.khan):

Now shuffle, x^2 +3x - 2x - 6 = x^2 + 5x + 2x + 10 x -7x = 10 + 6 -6x = 16 x = -16/6

OpenStudy (saifoo.khan):

Finally, x = -8/3

OpenStudy (anonymous):

thank you everyone.